# Equidistribution of polynomially bounded o-minimal curves in homogeneous spaces

### v1

@@ 3. $(C,\alpha)$-good property
Let $\lVert\cdot\rVert_{1}$ denote the $\ell_{1}$-norm on $\mathbb{R}^{N}$, which is the sum of absolute values of the coordinates. Let
(3.11) $\hat{\underline{c}}(s):=\frac{\underline{c}(s)}{\|\underline{c}(s)\|_{1}},\,\forall s\geq 1.$
We observe that by o-minimality, since $\lVert\hat{\underline{c}}(s)\rVert_{1}=1$, $\hat{\underline{c}}(s)$ converges to some vector v with $\|\textbf{v}\|_{1}=1$. Since $\underline{c}(s)\in\tilde{\mathscr{F}}\setminus\{0\}$ and $\tilde{\mathscr{F}}$ is a closed cone, we get $v\in\tilde{\mathscr{F}}\setminus\{0\}$.
Case 1: $\kappa\leq 0$. In this case, we treat two sub-cases in which $l(s)$ is bounded or not.
Case 1.1. $l(s)$ is bounded in $s$. Then $a(s)$ is also bounded, and by o-minimality $l(s)$ and $a(s)$ converge as $s\to\infty$. We first observe that $\lim_{s\to\infty}l(s)\neq 0$. In fact, if otherwise $\lim_{s\to\infty}l(s)=0$, then $\lim_{s\to\infty}\frac{a(s)}{l(s)}=\infty$ since $a(s)\geq T_{0}\geq 1,\forall s\geq 1$, which is a contradiction to the assumption that $\kappa\leq 0$. Also, the end-points of the intervals $[\alpha(s),\alpha(s)+\delta l(s)]\subset[a(s),a(s)+l(s)]$ converge to the end-points of intervals $I_{\delta}\subset I\subseteq[T_{0},\infty)$ of positive length, where $\frac{|I_{\delta}|}{|I|}=\delta$. In particular, $\|f_{\hat{\underline{c}}(s)}\|_{[a(s),a(s)+l(s)]}$ is uniformly bounded in $s$. Now
@@
$=\frac{\|f_{\underline{c}(s)}\|_{[a(s),a(s)+l(s)]}}{\|f_{\underline{c}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}>s.$
Namely, we have
@@
Case 1.2. $l(s)$ is unbounded in $s$. Then, by o-minimality $l(s)\to\infty$.
For $s\geq 1$, we define
(3.12) $C^{+}(s)=\sum_{i=0}^{m}|c^{+}_{i}(s)|l(s)^{r_{i}},\,C^{-}(s)=\sum_{j=1}^{n}|c^{-}_{j}(s)|\text{, and }C(s)=C^{+}(s)+C^{-}(s).$
Then, the by properties of bounded definable functions, the following limit exists in $\mathbb{R}^{m+1}$.
(3.13) $(w^{+}_{0},...,w^{+}_{m})=\lim_{s\to\infty}\frac{1}{C(s)}(c^{+}_{0}(s)l(s)^{r_{0}},...,c^{+}_{m}(s)l(s)^{r_{m}}).$
We denote
$\varphi_{{\underline{c}}(s)}(t):=\frac{f_{\underline{c}(s)}(a(s)+tl(s))}{\lambda(s)},\,\forall t\in[0,1].$
We first want to verify the following claim to be used in further arguments.
Claim 1.
(3.14) $\sup_{s\geq S_{0}}\lVert\varphi_{{\underline{c}}(s)}\rVert_{[0,1]}<\infty.$
Moreover, for any $t_{0}\in(0,1)$.
@@
Proof of Claim 1.
For each $0\leq i\leq m$, there exists $\nu_{i}>0$ such that for all $t\in[0,1]$, we have
@@
Since $\frac{a(s)}{l(s)}$ is bounded, we have $\lim_{s\to\infty}\frac{a(s)}{l(s)}=x_{0}\in[0,\infty)$. We now make the following observations:
  • We can pick $S_{0}\geq 1$ such that $l(s)\geq 1$ for all $s\geq S_{0}$ and the following two statements hold.
@@
      $\sup_{t\in[0,1]}\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}}$ is bounded for all $s\geq S_{0}$ and for all $0\leq i\leq m$.
      This is immediate since $r_{i}\geq 0$, $\lim_{s\to\infty}\frac{a(s)}{l(s)}=x_{0}$ and $t\in[0,1]$.
@@
      $\sup_{t\in[0,1]}l(s)^{-\nu_{i}}\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}-\nu_{i}}$ is bounded for all $s\geq S_{0}$ and for all $0\leq i\leq m$.
      This is clear if $r_{i}-\nu_{i}\geq 0$. Now suppose that $r_{i}-\nu_{i}<0$. Then, for all $t\in[0,1]$, since $a(s)\geq T_{0}\geq 1$ and $l(s)\geq 1$,
      $l(s)^{-\nu_{i}}\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}-\nu_{i}}\leq l(s)^{-\nu_{i}}\left(\frac{a(s)}{l(s)}\right)^{r_{i}-\nu_{i}}=\frac{a(s)^{r_{i}-\nu_{i}}}{l(s)^{r_{i}}}\leq 1.$
  • Since, by definability, $\lvert a(s)/l(s)-x_{0}\rvert=O(l(s)^{-\epsilon})$ for some $\epsilon>0$, for each $0\leq i\leq m$, there exists $\epsilon_{i}>0$ such that as $s\to\infty$
    $\sup_{t\in[0,1]}\left|\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}}-\left(x_{0}+t\right)^{r_{i}}\right|=O(l(s)^{\epsilon_{i}}).$
  • For any $0\leq i\leq m$, as $s\to\infty$,
    $\sup_{t\in[0,1]}l(s)^{-\nu_{i}}\left|\frac{a(s)}{l(s)}+t\right|^{r_{i}-\nu_{i}}=O(l(s)^{-\nu_{i}}).$
  • Let $\nu:=\min\{\epsilon_{0},\ldots,\epsilon_{m},\nu_{0},\ldots,\nu_{n}\}>0$. Then, for all $0\leq i\leq m$,
    $\sup_{t\in[0,1]}\left|\frac{F^{+}_{i}(a(s)+tl(s))}{l(s)^{r_{i}}}-(x_{0}+t)^{r_{i}}\right|=O(l(s)^{-\nu}),$
Thus, as $s\to\infty$, uniformly for $t\in[0,1]$,
(3.17) $f^{+}_{\underline{c}(s)}(a(s)+tl(s))=\sum_{i=0}^{m}(c_{i}^{+}(s)l(s)^{r_{i}})\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}}+O(C^{+}(s)l(s)^{-\nu}).$
We also observe that as $s\to\infty$, we have
(3.18) $\sup_{t\in[0,1]}\left|f_{\underline{c}(s)}^{-}(a(s)+tl(s))\right|=O(C^{-}(s)),$
and for any $t_{0}\in(0,1)$,
(3.19) $\sup_{t\in[t_{0},1]}\left|f_{\underline{c}(s)}^{-}(a(s)+tl(s))\right|=O(C^{-}(s)l(s)^{-\kappa_{n}}).$
Now (3.14) follows from (3.17) and (3.18), and (3.15) follows from (3.17) and (3.19). This completes the proof of the Claim 1.
By (3.9) and the definition of the choice function $\phi$,
(3.20) $s<\frac{\|f_{\underline{c}(s)}\|_{[a(s),a(s)+l(s)]}}{\|f_{\underline{c}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}=\frac{\|\varphi_{{\underline{c}}(s)}\|_{[0,1]}}{\|\varphi_{{\underline{c}}(s)}\|_{I_{\delta}(s)}}$
where $I_{\delta}(s):=\left[\frac{\alpha(s)-a(s)}{l(s)},\frac{\alpha(s)-a(s)+\delta l(s)}{l(s)}\right]\subseteq[0,1]$ having length $|I_{\delta}(s)|=\delta$ for all $s\geq 1.$ Therefore, by (3.14),
(3.21) $\lim_{s\to\infty}\|\varphi_{{\underline{c}}(s)}\|_{I_{\delta}(s)}\leq\lim_{s\to\infty}\frac{\|\varphi_{{\underline{c}}(s)}\|_{[0,1]}}{s}=0.$
And, since the end-points of $I_{\delta}(s)$ are bounded, they converge to endpoints of a sub-interval $I_{\delta}$ of $[0,1]$ of length $\delta$. Therefore, from (3.15) and (3.21), we deduce that
$w^{+}_{0}\left(x_{0}+t\right)^{r_{0}}+\cdots+w^{+}_{m}\left(x_{0}+t\right)^{r_{m}}=0,$
for all $t\in I_{\delta}\cap[t_{0},1]$ for any $t_{0}\in(0,1)$; and hence for all $t\in I_{\delta}$. Since $I_{\delta}$ is an interval of length $\delta$, by Lemma 3.5, we obtain that $w^{+}_{0}=\cdots=w^{+}_{m}=0$. Therefore, by (3.13), we get
@@
In view of (3.11), we have
@@
$=\lim_{s\to\infty}\frac{(c_{1}^{-}(s),...,c_{n}^{-}(s),c^{+}_{0}(s),...,c^{+}_{m}(s))}{\|(c^{-}_{1}(s),...,c^{-}_{n}(s),c^{+}_{0}(s),...,c^{+}_{m}(s))\|_{1}}.$
Then, $\|\textbf{v}\|_{1}=1$ and $\textbf{v}\in\tilde{\mathscr{F}}\smallsetminus\{0\}$.
By (3.22), we have
@@
$\leq\lim_{s\to\infty}\frac{\|(c_{1}^{-}(s),...,c_{n}^{-}(s))\|_{1}}{\|(c^{-}_{1}(s),...,c^{-}_{n}(s),c^{+}_{0}(s),...,c^{+}_{m}(s))\|_{1}}\text{, as $l(s)\to\infty$ and $r_{i}\geq 0$}$
@@
Therefore, $\|(v_{1}^{-},...,v_{n}^{-})\|_{1}$. As a consequence, we get $v^{+}_{0}=\cdots=v^{+}_{m}=0$. Therefore, $f_{\textbf{v}}(t)\to 0$ as $t\to\infty$, where $\textbf{v}\in\tilde{\mathscr{F}}\smallsetminus\{0\}$, and hence $f_{\textbf{v}}\in\mathscr{F}\smallsetminus\{0\}$. This contradicts our assumption that no nonzero function in $\mathscr{F}$ decays to zero.
Remark 3.9.
This is the only place in the proof of Theorem 1.14 where we use the non-decaying function assumption for the closed cone $\mathscr{F}$. And this is the main reason we need the non-contraction assumption on the curves in this article.
Case 2: $\kappa>0$.
Let $\lambda(s)=\|\underline{c}(s)\|_{1}$ and denote: $\hat{\underline{c}}:=\frac{\underline{c}}{\lambda}$. As above, note that
$\frac{\|f_{\hat{\underline{c}}(s)}\|_{[a(s),a(s)+l(s)]}}{\|f_{\hat{\underline{c}}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}=\frac{\|f_{\underline{c}(s)}\|_{[a(s),a(s)+l(s)]}}{\|f_{\underline{c}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}>s.$
We now consider two cases: $a(s)$ is bounded or $a(s)\to\infty$.
Case 2.1. $a(s)$ is bounded. We observe that $\lim_{s\to\infty}l(s)=0$. This follows since $a(s)\geq T_{0}\geq 1$ and $\kappa>0$, we get $\frac{a(s)}{l(s)}\sim C_{0}s^{\kappa}\to\infty$. We have that $\lim_{s\to\infty}a(s)=x_{0}\geq T_{0}$, and $\lim_{s\to\infty}\hat{\underline{c}}(s)=\textbf{v}$ with $\|\textbf{v}\|_{1}=1.$
By our choice of $T_{0}$, the basis functions $F^{-}_{i},F^{+}_{j}$ are $(N+1)$-times continuously differentiable where $N:=(n+m+1)$. For any $t\geq T_{0}$, let $W(t)$ denote the transpose of the Wronskian matrix $W(F_{-n}(t),\ldots,F_{m}(t))$, whose $k$-th row is $(\partial^{k}F_{-n}(t),\ldots,\partial^{k}F_{m}(t))$ for $k=0,\dots,N$. For $h\in\mathbb{R}$, define
(3.23) $Q_{N,a(s)}(h)=\hat{\underline{c}}(s)W(a(s))\cdot(h^{0},\ldots,h^{N}),$
where $\cdot$ denotes the dot product on $\mathbb{R}^{N+1}$. By Taylor’s theorem (mean value form for the remainder), for all $h\in[0,l(s)]$, there exists $\xi\in a(s)+[0,l(s)]:=[a(s),a(s)+l(s)]$ such that
@@
Since $a(s)$, $\hat{\underline{c}}(s)$, and $l(s)$ are bounded, we get that $|f_{\hat{\underline{c}}(s)}^{(N+1)}(\xi)|$ is uniformly bounded in the range $\xi\in[a(s),a(s)+l(s)]$.
Denote $I(s):=[0,l(s)]$, and $I_{\delta}(s):=[0,\delta l(s)]$. We have:
(By Taylor approximation) $\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I(s)}=$ $\|Q_{N,a(s)}\|_{I(s)}+O(l(s)^{N+1})$
(By Proposition 3.4) $\leq$ $M_{N}(\delta)\|Q_{N,a(s)}\|_{I_{\delta}(s)}+O(l(s)^{N+1})$
(By Taylor approximation) $\leq$ $M_{N}(\delta)\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}+O(l(s)^{N+1})$
@@
Since $\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}=\|Q_{N,a(s)}\|_{I_{\delta}(s)}+O(l(s)^{N+1})$, we will show that
@@
This outcome gives a contradiction to our assumption that $s\leq\frac{\|f_{{\underline{c}}(s)}\|_{I(s)}}{\|f_{{\underline{c}}(s)}\|_{I_{\delta}(s)}}$ for all $s\geq 1$.
To prove (3.24), we observe the following if $q(x)=c_{0}+c_{1}x+...+c_{k}x^{k}$ is a polynomial of degree bounded by $k$ and $J\subset[0,1]$ is an interval of length $|J|<1$, then by Proposition 3.4,
(3.25) $\|q\|_{J}\geq(k+1)^{-1}k^{-k}\|q\|_{[0,1]}|J|^{k}\geq\|(c_{0},c_{1},...,c_{k})\|_{1}|J|^{k},$
where the implied constant depends only on $k$, because $\|q\|_{[0,1]}$ and $\|\underline{c}\|_{1}$ are two norms on the $(k+1)$-dimensional space of polynomials of degree at most $k$.
By (3.23), $Q_{N,a(s)}$ is a polynomial of degree bounded by $N$ whose coefficients are given by $W(a(s))\hat{\underline{c}}(s)$. We have $W(a(s))\to W(x_{0})$ as $s\to\infty$. And, by Lemma 3.6, $W(x_{0})$ is nonsingular. Therefore,
$\lVert W(a(s))\hat{\underline{c}}(s)\rVert_{1}\geq\lVert W(a(s))^{-1}\rVert_{1}^{-1}\lVert\hat{\underline{c}}(s)\rVert_{1}\geq(1/2)\lVert W(x_{0})^{-1}\rVert_{1}^{-1}>0$
for all $s\gg 1$, as $\lVert\hat{\underline{c}}(s)\rVert_{1}=1$. Then, by (3.25) applied to $J=[0,\delta l(s)]$, we get
(3.26) $\|Q_{N,a(s)}\|_{I_{\delta}(s)}\gg\lVert W(x_{0})^{-1}\rVert_{1}^{-1}\delta^{N}l(s)^{N}.$

### v2

@@ 3. $(C,\alpha)$-good property
The remainder of the proof is organized by dividing the argument into cases based on the sign of $\kappa$ and on whether $l(s)$ and $a(s)$ are bounded or unbounded. In each case, we derive a contradiction. To guide the reader, we begin with a brief overview.
If $\kappa\leq 0$, this implies that the length of the interval $I(s)=[a(s),a(s)+l(s)]$ grows faster than its starting point. The case where $l(s)$ is bounded is easier to handle, since in that case $I(s)$ remains a bounded interval. When $l(s)$ is unbounded, we exploit the fact that polynomially bounded definable functions $f(t)$ are asymptotically equivalent to functions of the form $ct^{r}$.
On the other hand, if $\kappa>0$, the intervals $I(s)$ are short relative to their left endpoints. In this case, we can apply Taylor approximation, and use the $\delta$-goodness of polynomials (see Proposition 3.5) to reach a contradiction.
Before proceeding to the case analysis, we record the following observation, which will be useful throughout the proof. Let $\|\cdot\|_{1}$ denote the $L^{1}$-norm on $\mathbb{R}^{N}$, and define
(3.8) $\hat{\underline{c}}(s):=\frac{\underline{c}(s)}{\|\underline{c}(s)\|_{1}},\ s\geq 1.$
The function $s\mapsto\hat{\underline{c}}(s)$ is a bounded definable function with $\|\hat{\underline{c}}(s)\|_{1}=1$. It follows that
$\lim_{s\to\infty}\hat{\underline{c}}(s)=\mathbf{v}$
for some $\mathbf{v}\in\mathbb{R}^{N}$ satisfying $\|\mathbf{v}\|_{1}=1.$ Since $\underline{c}(s)\in\tilde{\mathscr{F}}\setminus\{0\}$ and $\tilde{\mathscr{F}}$ is a closed cone, we conclude that $\mathbf{v}\in\tilde{\mathscr{F}}\setminus\{0\}$.
Case 1: $\kappa\leq 0$. There are two sub-cases: $\{l(s):s\geq 1\}$ is bounded or not.
Case 1.1. $l(s)$ is bounded in $s$. Then $a(s)$ is also bounded, and by o-minimality $l(s)$ and $a(s)$ converge as $s\to\infty$. We observe that
$\lim_{s\to\infty}l(s)\neq 0.$
Because, since $a(s)\geq T_{0}\geq 1,\forall s\geq 1$, if $\lim_{s\to\infty}l(s)=0$, then $\lim_{s\to\infty}\frac{a(s)}{l(s)}=\infty$, which contradicts the assumption that $\kappa\leq 0$. We also note that the end-points of the nested intervals
$[\alpha(s),\alpha(s)+\delta l(s)]\subseteq[a(s),a(s)+l(s)],$
converge to the end-points of intervals $I_{\delta}\subseteq I\subseteq[T_{0},\infty)$ of positive length, where $\frac{|I_{\delta}|}{|I|}=\delta$. Hence $\|f_{\hat{\underline{c}}(s)}\|_{[a(s),a(s)+l(s)]}$ is uniformly bounded in $s$. Now,
@@
$=\frac{\|f_{\underline{c}(s)}\|_{[a(s),a(s)+l(s)]}}{\|f_{\underline{c}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}>s,$
so that
@@
Case 1.2. $l(s)$ is unbounded in $s$. In this case, we will achieve a contradiction by proving that:
$\lim_{s\to\infty}\hat{}\underline{c}(s)=(v_{1}^{-},...,v_{n}^{-},0,...,0),$
where $(v_{1}^{-},...,v_{n}^{-})\neq 0.$ This leads to a contradiction, because
$f_{(v_{1}^{-},...,v_{n}^{-},0,...,0)}\in\mathscr{F}\setminus\{0\}\text{ and }\lim_{t\to\infty}f_{(v_{1}^{-},...,v_{n}^{-},0,...,0)}(t)=0,$
is contrary to our assumption on $\mathscr{F}$.
Remark 3.11.
This is the only place in the proof of Theorem 1.12 where we use the non-decaying function assumption for the closed cone $\mathscr{F}$, and it highlights the necessity of the non-contracting assumption for obtaining the $C,\alpha$-good property.
We now proceed with our proof. Since $l(s)$ is an unbounded definable function, we have that $l(s)\to\infty$. For $s\geq 1$, we define
(3.9) $C^{+}(s):=\sum_{i=0}^{m}|c^{+}_{i}(s)|l(s)^{r_{i}},\,C^{-}(s):=\sum_{j=1}^{n}|c^{-}_{j}(s)|\text{, and }C(s):=C^{+}(s)+C^{-}(s).$
Since $s\mapsto\frac{(c^{+}_{0}(s)l(s)^{r_{0}},...,c^{+}_{m}(s)l(s)^{r_{m}})}{C(s)}$ is definable and bounded, we get that:
(3.10) $(w^{+}_{0},...,w^{+}_{m}):=\lim_{s\to\infty}\frac{(c^{+}_{0}(s)l(s)^{r_{0}},...,c^{+}_{m}(s)l(s)^{r_{m}})}{C(s)}\in[0,1]^{m+1}.$
Claim.
There exists an interval $J\subset[0,1]$ of positive length such that
$w_{0}^{+}t^{r_{0}}+\cdots+w_{m}^{+}t^{r_{m}}=0,~{}\forall t\in J.$
Here, $r_{i}$ are as in (3.5). In particular, $(w^{+}_{0},...,w^{+}_{m})=0$.
Proof of Claim.
We consider
$\varphi_{{\underline{c}}(s)}(t):=\frac{f_{\underline{c}(s)}(a(s)+tl(s))}{C(s)},\ \forall t\in[0,1].$
Recall that, by our choice function for the definable set (3.6),
(3.11) $s<\frac{\|f_{\underline{c}(s)}\|_{[a(s),a(s)+l(s)]}}{\|f_{\underline{c}(s)}\|_{[\alpha(s),\alpha(s)+\delta l(s)]}}=\frac{\|\varphi_{{\underline{c}}(s)}\|_{[0,1]}}{\|\varphi_{{\underline{c}}(s)}\|_{I_{\delta}(s)}},$
where $I_{\delta}(s):=\left[\frac{\alpha(s)-a(s)}{l(s)},\frac{\alpha(s)-a(s)+\delta l(s)}{l(s)}\right]\subseteq[0,1]$ having length $|I_{\delta}(s)|=\delta$ for all $s\geq 1.$ Since $\frac{a(s)}{l(s)}$ is a bounded definable function, we have
$\lim_{s\to\infty}\frac{a(s)}{l(s)}=x_{0}\in[0,\infty).$
Also, since the end-points of $I_{\delta}(s)$ are bounded, they converge to end-points of a sub-interval $I_{\delta}$ of $[0,1]$ of length $\delta$. We will prove that:
(3.12) $\limsup_{s\to\infty}\lVert\varphi_{{\underline{c}}(s)}\rVert_{I_{\delta}(s)}=0,$
and that for any $t_{0}\in(0,1)$:
@@
We note that the claim readily follows from (3.12) and (3.13). To prove (3.12), it is sufficient to show that
$\sup_{s\to\infty}\|\varphi_{\underline{c}(s)}\|_{[0,1]}<\infty,$
see (3.11). For each $0\leq i\leq m$, there exists a $\nu_{i}>0$ such that for all $t\in[0,1]$, we have as $l\to\infty$,
@@
We can pick $S_{0}\geq 1$ such that $l(s)\geq 1$ for all $s\geq S_{0}$ and that the following two statements hold:
@@
    $\sup_{t\in[0,1]}\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}}$ is uniformly bounded in $s\geq S_{0}$ and $0\leq i\leq m$.
    Indeed, because $r_{i}\geq 0$, $\lim_{s\to\infty}\frac{a(s)}{l(s)}=x_{0}$, and $t\in[0,1]$.
@@
    $\sup_{t\in[0,1]}l(s)^{-\nu_{i}}\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}-\nu_{i}}$ is uniformly bounded in $s\geq S_{0}$ and $0\leq i\leq m$.
    This is clear if $r_{i}-\nu_{i}\geq 0$. Now suppose that $r_{i}-\nu_{i}<0$. Then, for all $t\in[0,1]$,
    $l(s)^{-\nu_{i}}\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}-\nu_{i}}\leq$ $l(s)^{-\nu_{i}}\left(\frac{a(s)}{l(s)}\right)^{r_{i}-\nu_{i}}$
    $=$ $\frac{a(s)^{r_{i}-\nu_{i}}}{l(s)^{r_{i}}}=O(1),$
    since $\frac{a(s)}{l(s)}=O(s^{\kappa})$ with $\kappa\leq 0$, and since $a(s)\geq T_{0}\geq 1$.
By (2), as $s\to\infty$, uniformly in $t\in[0,1]$,
(3.15) $f^{+}_{\underline{c}(s)}(a(s)+tl(s))=\sum_{i=0}^{m}(c_{i}^{+}(s)l(s)^{r_{i}})\left(\frac{a(s)}{l(s)}+t\right)^{r_{i}}+O(C^{+}(s)).$
In addition, we observe that as $s\to\infty$:
(3.16) $\sup_{t\in[0,1]}\left|f_{\underline{c}(s)}^{-}(a(s)+tl(s))\right|=O(C^{-}(s)).$
Here, $C^{+},C^{-}$ are given by (3.9). Now:
$|\varphi_{{\underline{c}}(s)}(t)|$ $=\frac{|f_{\underline{c}(s)}(a(s)+tl(s))|}{C(s)}$
$=\frac{|f^{+}_{\underline{c}(s)}(a(s)+tl(s))+f_{\underline{c}(s)}^{-}(a(s)+tl(s))|}{C(s)}=O(1),~{}\text{as $s\to\infty$,}$
uniformly in $t\in[0,1]$. This proves (3.12).
Next, we prove (3.13). We have $\lim_{s\to\infty}l(s)=\infty$, $a(s)/l(s)>0$, and $\lim_{s\to\infty}a(s)/l(s)=x_{0}$. Fix $t_{0}\in(0,1)$. Then $t_{0}\leq\lvert a(s)/l(s)+t\rvert\leq x_{0}+2$ for all $t\in[t_{0},1]$, for all large $s$. So, for any $0\leq i\leq m$, by (3.14), as $s\to\infty$,
$\sup_{t\in[t_{0},1]}\left|\frac{F^{+}_{i}(a(s)+tl(s))}{l(s)^{r_{i}}}-(x_{0}+t)^{r_{i}}\right|=O(l(s)^{-\nu_{i}}).$
Let $\nu=\min\{\nu_{0},\ldots,\nu_{m}\}>0$. Then, as $s\to\infty$, uniformly in $t\in[t_{0},1]$,
(3.17) $f^{+}_{\underline{c}(s)}(a(s)+tl(s))=\sum_{i=0}^{m}(c_{i}^{+}(s)l(s)^{r_{i}})(x_{0}+t)^{r_{i}}+O(C^{+}(s)l(s)^{-\nu}).$
Also,
(3.18) $\sup_{t\in[t_{0},1]}\left|f_{\underline{c}(s)}^{-}(a(s)+tl(s))\right|=O(C^{-}(s)l(s)^{-\kappa_{n}}),$
where $\kappa_{n}$ is as in (3.5). By (3.17) and (3.18), for $t\in[t_{0},1]$, we have
||| $\varphi_{\underline{c}(s)}(t)-(w^{+}_{0}(x_{0}+t)^{r_{0}}+\cdots+w^{+}_{m}(x_{0}+t)^{r_{m}})|$
$\leq\sum_{i=0}^{m}\left|\frac{c_{i}^{+}(s)l(s)^{r_{i}}}{C(s)}-w_{i}^{+}\right|(x_{0}+t)^{r_{i}}+O(l(s)^{-\kappa_{n}}+l(s)^{-\nu}),$
and the last expression decays to 00 uniformly for $t\in[t_{0},1]$, as $s\to\infty$. ∎
Using our claim and by recalling (3.10), we get that
$\lim_{s\to\infty}\frac{(c_{1}^{-}(s),...,c_{n}^{-}(s),c^{+}_{0}(s)l(s)^{r_{0}},...,c^{+}_{m}(s)l(s)^{r_{m}})}{C(s)}$
@@
We now again consider ^$\hat{}\underline{c}(s)$, as in (3.8). We have
@@
$=\lim_{s\to\infty}\frac{(c_{1}^{-}(s),...,c_{n}^{-}(s),c^{+}_{0}(s),...,c^{+}_{m}(s))}{\|(c^{-}_{1}(s),...,c^{-}_{n}(s),c^{+}_{0}(s),...,c^{+}_{m}(s))\|_{1}},$
where $\|\textbf{v}\|_{1}=1$ and $\textbf{v}\in\tilde{\mathscr{F}}\setminus\{0\}$. By (3.19), and since $l(s)\to\infty$ and $r_{i}\geq 0$:
@@
$\leq\lim_{s\to\infty}\frac{\|(c_{1}^{-}(s),...,c_{n}^{-}(s))\|_{1}}{\|(c^{-}_{1}(s),...,c^{-}_{n}(s),c^{+}_{0}(s),...,c^{+}_{m}(s))\|_{1}}$
@@
Therefore, $\|(v_{1}^{-},...,v_{n}^{-})\|_{1}=1$. As a consequence, we get $v^{+}_{0}=\cdots=v^{+}_{m}=0.$
Case 2: $\kappa>0$. Here, we consider two cases corresponding to $a(s)$ being bounded or $a(s)\to\infty$.
Case 2.1. $a(s)$ is bounded. We have that $\lim_{s\to\infty}a(s)=x_{0}\geq T_{0}$. Since $\lim_{s\to\infty}\frac{a(s)}{l(s)}=\infty$, it follows that $\lim_{s\to\infty}l(s)=0$. We again consider ^$\hat{}\underline{c}(s)$, as in (3.8), and recall $\lim_{s\to\infty}\hat{\underline{c}}(s)=\textbf{v}$ with $\textbf{v}\in\tilde{\mathscr{F}}$ such that $\|\textbf{v}\|_{1}=1.$
By our choice of $T_{0}$, the basis functions $F^{-}_{i},F^{+}_{j}$ are $(N+1)$-times continuously differentiable, where $N:=n+m+1$. For $t\geq T_{0}$, let $W(t)$ denote the Wronskian matrix $W(F_{-n}(t),\ldots,F_{m}(t))$, whose $k$-th row is $(F_{-n}^{(k)}(t),\ldots,F_{m}^{(k)}(t))$, for $k=0,\dots,N$. The Taylor polynomial of order $N$ of the function $h\mapsto f_{\hat{\underline{c}(s)}}(a(s)+h)$ centered at $h=0$ is expressed by:
(3.20) $Q_{N,a(s)}(h)=\hat{\underline{c}}(s)W(a(s))^{\top}\cdot(h^{0},\ldots,h^{N}),$
where $\cdot$ denotes the dot product on $\mathbb{R}^{N+1}$. By Taylor’s theorem (mean value form for the remainder), for all $h\in[0,l(s)]$, there exists a $\xi\in a(s)+[0,l(s)]$ such that
@@
Since $a(s)$, $\hat{\underline{c}}(s)$, and $l(s)$ are bounded, we obtain that $|f_{\hat{\underline{c}}(s)}^{(N+1)}(\xi)|$ is uniformly bounded.
Denote $I(s):=[0,l(s)]$, and let $I_{\delta}(s):=[0,\delta l(s)]$. Thus, we conclude:
$\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I(s)}=\|Q_{N,a(s)}\|_{I(s)}+O(l(s)^{N+1}),$
and similarly,
(3.21) $\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}=\|Q_{N,a(s)}\|_{I_{\delta}(s)}+O(l(s)^{N+1}).$
Using Proposition 3.5,
$\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I(s)}=$ $\|Q_{N,a(s)}\|_{I(s)}+O(l(s)^{N+1})$
$\leq$ $M_{N}(\delta)\|Q_{N,a(s)}\|_{I_{\delta}(s)}+O(l(s)^{N+1})$
$\leq$ $M_{N}(\delta)\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}+O(l(s)^{N+1})$
@@
We will now show that
@@
Before proceeding with the proof, note that by combining (3.22) with (3.21), we obtain that $\frac{l(s)^{N+1}}{\|f_{\hat{\underline{c}}(s)}\|_{a(s)+I_{\delta}(s)}}$ is bounded for all large $s$. This is a contradiction, since we are assuming that $s\leq\frac{\|f_{{\underline{c}}(s)}\|_{I(s)}}{\|f_{{\underline{c}}(s)}\|_{I_{\delta}(s)}}$ for all $s\geq 1$.
We proceed to prove (3.22). Observe that for a polynomial $q(x)=c_{0}+c_{1}x+...+c_{k}x^{k}$ of degree bounded by $k$ and an interval $J\subseteq[0,1]$, it follows by Proposition 3.5,
(3.23) $\|q\|_{J}\geq(k+1)^{-1}k^{-k}\|q\|_{[0,1]}|J|^{k}\gg\|(c_{0},c_{1},...,c_{k})\|_{1}|J|^{k}.$
Here, we have also used that $\|q\|_{[0,1]}$ and $\|\underline{c}\|_{1}$ are two norms on the $(k+1)$-dimensional space of polynomials of degree at most $k$, and we note that the implied constant depends only on $k$. By (3.20), $Q_{N,a(s)}$ is a polynomial of degree bounded by $N$ whose coefficients are given by $\hat{\underline{c}}(s)W(a(s))^{\top}$. We have that $W(a(s))^{\top}\to W(x_{0})^{\top}$ as $s\to\infty$, and we recall Lemma 3.8 which implies that $W(x_{0})^{\top}$ is non-singular. Therefore,
$\lVert\hat{\underline{c}}(s)W(a(s))^{\top}\rVert\gg\lVert\hat{\underline{c}}(s)\rVert_{1}\lVert(W(a(s))^{\top})^{-1}\rVert^{-1}=\lVert(W(x_{0})^{\top})^{-1}\rVert^{-1}>0$
for all $s\gg 1$. Then, by (3.23) applied to $J=[0,\delta l(s)]$, we get
(3.24) $\|Q_{N,a(s)}\|_{I_{\delta}(s)}\gg\lVert(W(x_{0})^{\top})^{-1}\rVert^{-1}\delta^{N}l(s)^{N}.$